# Extrinsic contribution to bosonic thermal Hall transport

### v1

@@ V.1 General bosonic model: physical picture
For the model without disorder, we start from a low-energy perspective and consider a generic nonlinear sigma-model for emergent bosonic $\bm{m},\bm{n}$ fields, conjugated with each other. We discuss two physical instances of such a field theory below. We assume, for simplicity, that the theory is rotationally invariant around the $z$ axis. Then the most general hamiltonian density preserving this symmetry, to the second order in the $\bm{m},\bm{n}$ fields and up to irrelevant terms 555(in the sense that they contain higher powers of spatial derivatives than another similar term already present in the hamiltonian density), is
$H$ $=\frac{1}{2\chi}\bm{m}^{2}-\tfrac{1}{2}c_{1}\bm{n}\cdot\nabla^{2}\bm{n}+\tfrac{1}{2}c_{2}(\nabla\bm{n})^{2}+\Delta\,\bm{n}^{2}$
@@
where $\chi,\Delta,c_{1},c_{2},d_{1},d_{2},d_{3},d_{4}$ are real parameters, and stability requires $\chi>0,c_{1}>0,c_{1}+c_{2}>0,\Delta\geq 0$).
Physical systems described by this hamiltonian include:
@@
Generalized elasticity – In this case, $H$ describes the low-energy lattice dynamics, with $\bm{n}$ the lattice local displacement field (usually called $\bm{u}$), $\bm{m}$ is the lattice local momentum field (usually called $\bm{p}$), $c_{1}$ and $c_{2}$ are elasticity moduli Landau and Lifshitz (1970), $\Delta=0$ is imposed by translational invariance (i.e. acoustic phonons are Goldstone bosons), $\chi^{-1}$ is the lattice mass density (usually called $\rho$), and $d_{i}\,(i=1..4)$ are phonon viscosity parameters Avron et al. (1995); Barkeshli et al. (2012). Such terms are known Qin et al. to generate a phonon Berry curvature and a phonon thermal Hall effect. Since these viscosity terms break time reversal symmetry, they may only arise in an effective phonon theory beyond pure elasticity, for instance from an external magnetic field beyond the Born-Oppenheimer approximation Saito et al. (2019) or as a result of spin-phonon interactions in magnetic materials Ye et al. .
Magnetic systems – In this case, $H$ describes the low-energy dynamics of spins. Here we discuss briefly how it can be derived from a microscopic theory, starting from a spin hamiltonian
$H_{\rm spin}$ $=\sum_{\mathbf{r},\mathbf{r}^{\prime}}J_{\mathbf{r},\mathbf{r}^{\prime}}^{ab}S_{\mathbf{r}}^{a}S_{\mathbf{r}^{\prime}}^{b}$ (44)
with $a,b\in\{x,y,z\}$. Assuming a bipartite lattice with Néel antiferromagnetic ordering along $\hat{\mathbf{z}}$, spin fluctuations can be described in terms of the local net magnetization $\bm{m}$ and local staggered magnetization fluctuations $\bm{n}$, which are two-component vectors orthogonal to the Néel order parameter and with each other. For a concrete illustration one can consider the square lattice and a Kitaev-Heisenberg magnetic exchange
$J_{\mathbf{r},\mathbf{r}^{\prime}}^{ab}=J\;{\mathbb{I}}_{\mathbf{r}^{\prime}\in{\rm N}(\mathbf{r})}\delta_{a,b}+K\;{\mathbb{I}}_{\mathbf{r}^{\prime}\in{\rm N}_{a}(\mathbf{r})}\delta_{a,b}$
@@
where ${\rm N}(\mathbf{r})$, respectively ${\rm N}_{a}(\mathbf{r})$, is the set of nearest neighbors of $\mathbf{r}$, respectively along a bond of type $a\in\{x,y\}$, and ${\mathbb{I}}_{\mathcal{E}}$ is the indicator of $\mathcal{E}$. To restore rotational symmetry in the effective theory, we select the rotation-invariant component of the hamiltonian, which then takes the form Eq.(V.1) where one identifies $\chi^{-1}=8J+2K,c_{1}={\mathtt{a}}^{2}(J+K/8),c_{2}={\mathtt{a}}^{2}K/4$, with ${\mathtt{a}}$ the lattice parameter (see details in Appendix G.4). The magnon mass $\Delta$ may result from anisotropic exchange, and non-vanishing $d_{i}$ coefficients require the effective breaking of time reversal symmetry in the magnon dynamics, e.g. from a ferromagnetic moment induced by Dzyaloshinskii-Moriya exchange or by Zeeman coupling to an exteral magnetic field. We do not consider this in more detail here, referring the reader to the literature on magnon Berry curvature McClarty (2022); Mook et al. (2017), and in the following we come back to the general hamiltonian Eq.(V.1) without reference to its origin from a microscopic spin model.
V.2 The model without disorder
@@
with convention $\epsilon_{12}=+$ and where $\eta=d_{1}-\tfrac{1}{2}d_{2}$. For details see Appendix G.1.
We note that $\eta$ and $d_{3}$ play essentially the same role in inducing a magnon Berry curvature, but with different momentum dependences. Because the $d_{3}$ term makes the Berry curvature very singular close to zero momentum, so that physical properties such as the thermal Hall conductivity may depend strongly on the gap $\Delta$, for convenience we assume in the following $d_{3}=0$ together with $\Delta=0$.
@@
${\sf\Gamma}$ $=\frac{i}{\chi}\begin{bmatrix}0&\delta_{i,j}\\ -\delta_{i,j}&-\eta\epsilon_{i,j}p^{2}\end{bmatrix},$ (49a)
${\sf H}$ $=\begin{bmatrix}p_{\mu}\left(c_{1}\delta_{\mu,\nu}\delta_{i,j}+c_{2}\delta_{i,\mu}\delta_{\nu,j}\right)p_{\nu}&0\\ 0&\chi\delta_{i,j}\end{bmatrix},$ (49b)
whence the dynamical matrix $\mathsf{K}=\mathsf{\Gamma}\mathsf{H}$. This defines the clean theory: we now investigate several instances of disorder.
V.3 Three models with disorder
Details of the calculations are provided in App. G.3, and results for $\kappa_{xy}^{\rm dis}$ are displayed in Fig. 1, with the numerical values given in Table 1.
@@ V.3 Three models with disorder
[…]
Table 1: Parameter values used for the numerical evaluations. Momentum integrals are computed over $\bm{p}\in[-\pi/{\mathtt{a}},+\pi/{\mathtt{a}}]^{2}$.
Case where $\mathsf{W}=\mathsf{K}$ For any clean theory $(\mathsf{H},\mathsf{\Gamma})$ with disorder $\hat{W}=\hat{\mathsf{H}}$ and thus $\mathsf{W}=\mathsf{K}$, corresponding physically to (smoothed) vacancies or a local modulation of the clean hamiltonian, interesting simplifications occur (see Appendix B.2) and the disorder-induced curvature identifies with the usual Berry curvature:
@@
×$\qquad\times\int\text{d}\varepsilon\,\delta(\varepsilon-{\sf K}_{\rm d})\,\varepsilon^{2}\partial_{\varepsilon}n_{\rm B}\left(\varepsilon,T\right),$ (51)
to be compared with Eq.(1). Crucially, this result is general in that it relies only on the constraint $\mathsf{W}=\mathsf{K}$ but not on the clean model $(\mathsf{H},\mathsf{\Gamma})$ chosen – in Fig. 1 we display the result Eq.(V.3) in the particular case of the clean theory Eqs.(49a),(49b).
@@
where $\tilde{\eta}$ can be interpreted as a local random fluctuation amplitude of $\eta$. Details of the calculations are provided in App. G.3.1: the third term in Eq.(32) vanishes, the first one is order $\eta+O(\eta^{3})$ and the second one is order $\eta^{-1}+O(\eta)$, singular in the limit $\eta\rightarrow 0$. As already mentioned in generality, the disorder-induced curvature does not depend on the strength of disorder: $\tilde{\eta}$ does not appear in the result.
@@
where the second term generates chirality, and the first term ensures regular behavior – see App. G.3.1 for details. Only the second term in $(\mathsf{\Omega}^{\mathsf{W}}_{p_{\mu}p_{\nu}})_{n}(p)$, Eq.(32), is not zero, and $\mathsf{\Omega}^{\mathsf{W}}\propto\tilde{\eta}p^{2}$. This causes a finite thermal Hall conductivity to arise, as shown in Fig. 1(c). We state again that this stems purely from the elastic scattering with impurities of magnons with zero (intrinsic) Berry curvature.
V.4 Results and discussion
@@
These three examples show that the side-jump contribution to $\kappa_{xy}$ gives an order one correction to the intrinsic part. Since both its amplitude and its sign depend on the specifics of the disorder hamiltonian, it can enhance or reduce $\kappa_{xy}$ and may even lead to sign changes. In addition, even in the case of a trivial (i.e. TR preserving) clean system, special forms of disorder like local viscosity terms can lead to a non-vanishing THE. In these examples, $\mathsf{\Omega}_{p_{x},p_{y}}$ and $\mathsf{\Omega}^{\mathsf{W}}_{p_{x},p_{y}}$ both are concentrated close to the gamma-point $k=0$. It might be interesting, as another direct application, to study cases, for instance hybridized magnon-phonon bands, where one or both of these quantities is largest at a finite momentum, e.g. around the avoided-crossing point.
That the contribution we derived is non-perturbative in the disorder strength (i.e. it scales as $n_{\rm imp}^{0}$) is an intriguing, yet well-known Nagaosa et al. property of coordinate shift effects in anomalous Hall transport, and may be given a very intuitive explanation. Whenever heat is carried by well-defined quasiparticles, one may write $\kappa\sim c\,v\,l$ with $c$ the specific heat, $v$ the velocity and $l$ the mean-free path. This holds for any contribution independently, so focussing on the side-jump Hall conductivity,
$\kappa_{xy}^{\rm sj}\sim c\,v_{\perp}^{\rm sj}\,l,$ (55)
where $v_{\perp}^{\rm sj}$ is now the side-jump transverse velocity, $v_{\perp}^{\rm sj}\propto n_{\rm imp}$. Here we have assumed that $l$ is governed by disorder scattering, $l^{-1}\propto n_{\rm imp}$, which entails that $\kappa_{xy}^{\rm sj}$ is independent of $n_{\rm imp}$.

### v2

@@ V.3.1 Introduction and realization in physical systems
We now consider a formal application to a less specific instance of a bosonic model, formulated in the equivalent language of effective field theory. For the model without disorder, we start from a low-energy perspective and consider a nonlinear sigma-model for emergent bosonic $\bm{m},\bm{n}$ fields, conjugated with each other. We discuss two physical instances of such a field theory below. We assume, for simplicity, that the theory is rotationally invariant around the $z$ axis. Then the most general hamiltonian density preserving this symmetry, to the second order in the $\bm{m},\bm{n}$ fields and up to irrelevant terms 888(in the sense that they contain higher powers of spatial derivatives than another similar term already present in the hamiltonian density), is
$H$ $=\frac{1}{2\chi}\bm{m}^{2}-\tfrac{1}{2}c_{1}\bm{n}\cdot\nabla^{2}\bm{n}+\tfrac{1}{2}c_{2}(\nabla\bm{n})^{2}+\tfrac{1}{2}\Delta\,\bm{n}^{2}$
@@
where $\chi,\Delta,c_{1},c_{2},d_{1},d_{2},d_{3},d_{4}$ are real parameters, and stability requires $\chi>0,c_{1}>0,c_{1}+c_{2}>0,\Delta\geq 0$). The coefficients $d_{i}\,(i=1..4)$, by coupling together conjugated fields $\bm{m},\bm{n}$, will enable the breaking of time reversal symmetry. Of course, in lattice systems with enforced crystal symmetries, not all of these terms will be allowed (and possibly other terms, beyond our quadratic low-energy hamiltonian, will turn out to be more relevant). The actual magnitude of these terms (and of the resulting THE) may also vary vastly from one physical system to the other. Physical systems described by this hamiltonian Eq.(V.3.1) include:
@@
Elasticity with viscosity terms – In this case, $H$ describes the low-energy lattice dynamics, with $\bm{n}$ the lattice local displacement field (usually called $\bm{u}$), $\bm{m}$ is the lattice local momentum field (usually called $\bm{p}$), $c_{1}$ and $c_{2}$ are elasticity moduli Landau and Lifshitz (1970), $\Delta=0$ is imposed by translational invariance (i.e. acoustic phonons are Goldstone bosons), $\chi^{-1}$ is the lattice mass density (usually called $\rho$), and $d_{i}\,(i=1..4)$ are phonon viscosity parameters Avron et al. (1995); Barkeshli et al. (2012). Such terms are known Qin et al. to generate a phonon Berry curvature and a phonon thermal Hall effect. Since these viscosity terms break time reversal symmetry, they may only arise in an effective phonon theory beyond pure elasticity, for instance from an external magnetic field beyond the Born-Oppenheimer approximation Saito et al. (2019) or as a result of spin-phonon interactions in magnetic materials Ye et al. (2021).
@@
Magnetic systems – In this case, $H$ describes the low-energy dynamics of spins. Here we discuss briefly how it can be derived from a microscopic theory, starting from a spin hamiltonian $H_{\rm spin}=\sum_{\mathbf{r},\mathbf{r}^{\prime}}J_{\mathbf{r},\mathbf{r}^{\prime}}^{ab}S_{\mathbf{r}}^{a}S_{\mathbf{r}^{\prime}}^{b}$ with $a,b\in\{x,y,z\}$. Assuming a bipartite lattice with Néel antiferromagnetic ordering along $\hat{\mathbf{z}}$, spin fluctuations can be described in terms of the local net magnetization $\bm{m}$ and local staggered magnetization fluctuations $\bm{n}$, which are two-component vectors orthogonal to the Néel order parameter and with each other. For a concrete illustration one can consider the square lattice and a Kitaev-Heisenberg magnetic exchange $J_{\bm{r},\bm{r}^{\prime}}^{ab}=J\;{\mathbb{I}}_{\bm{r}^{\prime}\in{\rm N}(\bm{r})}\delta_{a,b}+K\;{\mathbb{I}}_{\bm{r}^{\prime}\in{\rm N}_{a}(\bm{r})}\delta_{a,b}$ where ${\rm N}(\bm{r})$, respectively ${\rm N}_{a}(\bm{r})$, is the set of nearest neighbors of $\bm{r}$, respectively along a bond of type $a\in\{x,y\}$, and $\mathbb{I}$$_{\mathcal{E}}$ is the indicator of $\mathcal{E}$. To restore rotational symmetry in the effective theory, we select the rotation-invariant component of the hamiltonian, which then takes the form Eq.(V.3.1) where one identifies $\chi^{-1}=8J+2K,c_{1}={\mathtt{a}}^{2}(J+K/8),c_{2}={\mathtt{a}}^{2}K/4$, with $\mathtt{a}$} the lattice parameter (see details in Appendix H.4). The magnon mass $\Delta$ may result from anisotropic exchange, and non-vanishing $d_{i}$ coefficients require the effective breaking of time reversal symmetry in the magnon dynamics, e.g. from a ferromagnetic moment induced by Dzyaloshinskii-Moriya exchange and Zeeman coupling to an exteral magnetic field. Although these coefficients are essential in the existence of a nonvanishing Hall effect, we do not consider their possible microscopic origin in more detail here, referring the reader to the literature on magnon Berry curvature McClarty (2022); Mook et al. (2017).
In the following we come back to the low-energy hamiltonian Eq.(V.3.1) without reference to its possible origin from a microscopic spin model.
V.3.2 The model without disorder
@@
with convention $\epsilon_{12}=+$ and where $\eta=d_{1}-\tfrac{1}{2}d_{2}$. For details see Appendix H.1.
@@
${\sf\Gamma}$ $=\frac{i}{\chi}\begin{bmatrix}0&\delta_{i,j}\\ -\delta_{i,j}&-\epsilon_{i,j}(\eta p^{2}-d_{3}/\chi)\end{bmatrix}$, (61a)
${\sf H}$ $=\begin{bmatrix}p_{\mu}\left(c_{1}\delta_{\mu,\nu}\delta_{i,j}+c_{2}\delta_{i,\mu}\delta_{\nu,j}\right)p_{\nu}+\Delta\delta_{i,j}&0\\ 0&\chi\delta_{i,j}\end{bmatrix}$, (61b)
whence the dynamical matrix $\mathsf{K}=\mathsf{\Gamma}\mathsf{H}$.
A boson Berry curvature, leading to an intrinsic Hall effect, is caused by the lower-right block of $\sf\Gamma$}, see Eq.(61a). We note that $\eta$ and $d_{3}$ will play essentially the same role in that regard, but with different momentum dependences. Because the $d_{3}$ term makes the Berry curvature very singular close to zero momentum, so that physical properties such as the (clean) thermal Hall conductivity may depend strongly on the gap $\Delta$, for convenience we assume in the following $d_{3}=0$.
These Eqs.(61a),(61b) define the clean theory: we now investigate several instances of disorder.
V.3.3 Three models with disorder
Details of the calculations are provided in App. H.3, and results for $\kappa_{xy}^{\rm dis}$ are displayed in Fig. 1, with the numerical values given in Table 1.
@@ V.3.3 Three models with disorder
$\mathtt{1.0}$} $\mathtt{1.0}$ $\mathtt{0.25}$ $\mathtt{-0.2}$ $\mathtt{+0.1}$ $\mathtt{0.0}$ $\mathtt{0.01}$ $\mathtt{0.2}$ $\mathtt{1}.0$
Table 1: Parameter values used for the numerical evaluations. Momentum integrals are computed over $\bm{p}\in[-\pi/{\mathtt{a}},+\pi/{\mathtt{a}}]^{2}$. We note that in the particular case where $d_{1}=0$, $d_{2}=0$, $d_{3}=0$ and $\tilde{\eta}=0$, TR symmetry is preserved and no THE is allowed.
Case where $\mathsf{W}=\mathsf{K}$ For any clean theory $(\mathsf{H},\mathsf{\Gamma})$ with disorder $\hat{W}=\hat{\mathsf{H}}$ and thus $\mathsf{W}=\mathsf{K}$, corresponding physically to locally modulating the clean hamiltonian by a random global prefactor, interesting simplifications occur (see Appendix B.2) and the disorder-induced curvature identifies with the usual Berry curvature:
@@
$\qquad\times\int\text{d}\varepsilon\,\delta(\varepsilon-{\sf K}_{\rm d})\,\varepsilon^{2}\partial_{\varepsilon}n_{\rm B}\left(\varepsilon,T\right),$ (63)
to be compared with Eq.(1). Crucially, this result is general in that it relies only on the constraint $\mathsf{W}=\mathsf{K}$ but not on the clean model $(\mathsf{H},\mathsf{\Gamma})$ chosen – in Fig. 1 we display the result Eq.(V.3.3) in the particular case of the clean theory Eqs.(61a),(61b).
@@
where $\tilde{\eta}$ can be interpreted as a local random fluctuation amplitude of $\eta$. Let us recall that in the clean case, $\eta$ is responsible for the breaking of TR symmetry and the generation of Berry curvature: in the disordered case, that role will be shared by $\tilde{\eta}$.
Details of the calculations are provided in App. H.3.1: the third term in Eq.(32) vanishes, the first one is order $\eta+O(\eta^{3})$ and the second one is order $\eta^{-1}+O(\eta)$, singular in the limit $\eta\rightarrow 0$. As already mentioned in generality, the disorder-induced curvature does not depend on the strength of disorder: $\tilde{\eta}$ does not appear in the final result.
@@
where only the second term generates chirality, and the first term ensures regular behavior – see App. H.3.1 for details. Only the second term in $(\mathsf{\Omega}^{\mathsf{W}}_{p_{\mu}p_{\nu}})_{n}(p)$, Eq.(32), is not zero, and $\mathsf{\Omega}^{\mathsf{W}}\propto\tilde{\eta}p^{2}$. This causes a finite thermal Hall conductivity to arise, as shown in Fig. 1(c). We state again that this stems purely from the elastic scattering with impurities of bosons with zero (intrinsic) Berry curvature.
V.3.4 Results and discussion
@@
These three examples show that the side-jump contribution to $\kappa_{xy}$ gives an order one correction to the intrinsic part. Since both its amplitude and its sign depend on the specifics of the disorder hamiltonian, it can enhance or reduce $\kappa_{xy}$ and may even lead to sign changes. In addition, even in the case of a trivial (i.e. TR preserving) clean system, special forms of disorder like local viscosity terms can lead to a non-vanishing THE.
The temperature dependences of $\kappa_{xy}^{\rm clean}$ and $\kappa_{xy}^{\rm dis}$ are imposed by the distribution in momentum space of $\mathsf{\Omega}_{p_{x}p_{y}}$ and $\mathsf{\Omega}_{p_{x}p_{y}}^{\mathsf{W}}$, respectively, and by the bosonic dispersion $\underline{\mathsf{K}}_{\rm d}$. An approximately linear dispersion $\underline{\mathsf{K}}_{\rm d}\sim k$ with a curvature going as $\mathsf{\Omega}^{(\mathsf{W})}\sim 1/k$ yield a thermal Hall conductivity $\kappa_{xy}\sim T^{2}$, as is seen in all three plots. The sharp peak in case (b) is governed by the interplay between the momentum distribution of $\mathsf{\Omega}^{\mathsf{W}}$ and the dispersion gap $\Delta$, so that its position is approximately $T\sim\Delta$. The maximum of the “broad peak” seen in all three plots corresponds to the saturation of the integral of the Berry curvature, which in our model is imposed by the Brillouin zone boundaries (i.e. the maximum occurs around $T\approx\theta_{\rm D}$ the Debye temperature of the boson bands). The orders of magnitude we provide here are purely qualitative, given that the quantitative determination of the Berry curvature (and similarly disorder-induced curvatures) of bosonic excitations in concrete models of materials is a challenge in its own right Saito et al. (2019).
Even these qualitative features are of course very model-dependent. In particular, in our examples, $\mathsf{\Omega}_{p_{x},p_{y}}$ and $\mathsf{\Omega}^{\mathsf{W}}_{p_{x},p_{y}}$ both are concentrated close to the gamma-point $k=0$, but it would be interesting, as another direct application, to study cases, for instance hybridized magnon-phonon bands, where one or both of these quantities is largest at a finite momentum, e.g. around the avoided-crossing point.
V.4 Experimental aspects
On the experimental side, the theory of side-jump developed above could help with the interpretation of thermal Hall measurements, especially in systems where heat is carried by chiral bosons such as spin excitations in quantum magnets. It is in principle possible to distinguish side-jump from other mechanisms of THE in cases where sample dependence for one given compound is observed, or by investigating dependence upon doping/substitutions. One key experimental feature of side-jump THE, due to the relation $\kappa_{xy}^{\rm sj}\propto l_{\rm net}/l_{\rm imp}$, is that $\kappa_{xy}$ as a function of $\kappa_{xx}$ is, as a rule of thumb, decreasing or constant, depending on whether $l_{\rm net}$ is dominated by interactions or disorder, respectively. This is at odds with skew-scattering mechanisms, where a thermal Hall resistivity $\rho_{xy}$ is generated from either disorder or interactions; then the thermal Hall conductivity $\kappa_{xy}\approx-\rho_{xy}/\rho_{xx}^{2}$ is, as a rule of thumb, increasing or constant as a function of $\kappa_{xx}\approx 1/\rho_{xx}$, depending on whether skew-scattering originates from interactions or disorder, respectively. This is also at odds with the intrinsic mechanism, where $\kappa_{xy}$ simply does not depend on $\kappa_{xx}$. These features are summarized in Table 2.
Mechanism generating THE Expected $\kappa_{xy}$ vs $\kappa_{xx}$
Intrinsic Constant
Skew-scattering (interactions) Increasing ($\kappa_{xy}\overset{\sim}{\propto}\kappa_{xx}^{2}$)
Skew-scattering (disorder) $\sim$ Constant
Side-jump (case $l_{\rm net}\ll l_{\rm imp}$) Decreasing ($\kappa_{xy}\overset{\sim}{\propto}\kappa_{xx}^{-1}$)
Side-jump (case $l_{\rm net}\approx l_{\rm imp}$) $\sim$ Constant
Table 2: Compared variations of $\kappa_{xy},\kappa_{xx}$ upon tuning disorder (expected behavior).
Experimental examples of compounds where $\kappa_{xy}$ is smaller for larger $\kappa_{xx}$ include the kagome disordered antiferromagnet Volborthite Watanabe et al. (2016), where the reported THE was attributed to spin excitations. Another example are pyrochlore terbium oxides, $\rm Tb_{2}Ti_{2}O_{7}$ Hirschberger et al. (2015a) and $\rm(Tb_{.3}Y_{.7})_{2}Ti_{2}O_{7}$ Hirokane et al. (2019), where upon substituting magnetic $\rm Tb^{3+}$ ions with nonmagnetic $\rm Y^{3+}$ one observes that $\kappa_{xx}$ slightly decreases while $\kappa_{xy}$ noticeably increases.  [^9] We note that the physics of thermal transport in fully nonmagnetic $\rm Y_{2}Ti_{2}O_{7}$ Sharma et al. (2024) appears to be quite different. While arguments have been put forward to discard either phonons or spin excitations as heat carriers in this family of compounds, we note that these observations are compatible with a non-negligible role of side-jump. A $\kappa_{xy}$ smaller for larger $\kappa_{xx}$ was also found Zhang et al. (2023) in the Kitaev compound $\alpha\text{-RuCl}_{3}$ – a fact that may be considered in regard with our Sec.V.2.
As a final note, while decreasing $\kappa_{xy}$ vs $\kappa_{xx}$ is in principle a sufficient argument supporting a side-jump mechanism, it is not necessary. Indeed, sample dependence with $\kappa_{xy},\kappa_{xx}$ roughly proportional to each other has been reported in kagome antiferromagnets Ca-kapellasite Doki et al. (2018) and Cd-kapellasite Akazawa et al. (2020) as well as in honeycomb ferromagnet $\rm VI_{3}$ Zhang et al. (2021), and in all three cases was attribued to spurious geometrical factors afflicting the data: unfortunately, such effects obscure possible signatures of side-jump mechanisms.